Imagine:

If a+b+c=4 and a^2+b^2+c^2=10 and then a^3+b^3+c^3=22 Then what does a^4+b^4+c^4=?

supie:

I think what we have to do is use this a+b+c=4 and a^2+b^2+c^2=10 and then a^3+b^3+c^3=22 To find what variable \[a,\ b,\ and\ c\ are\] Am I correct @Imagine?

Timmyspu:

u r correct @supie

Ruthy:

The and is 58 bcoz the distance between the first and to the second is 6 the second 17 then d last will be 28 making it 58

Imagine:

yes

Imagine:

@supie

supie:

Do you know the answer or is this a challenge?

supie:

I mean do you need the answer. or is it a challenge

Imagine:

I don't understand, how to add all them o get the answer.

supie:

Oh ok.

supie:

@jhonyy9

Imagine:

.

Imagine:

.-.

VAEHVAEH:

can u please close this because im tying to do hw and this is drawing my attention away from my work

Imagine:

this is my question? and you want me to close it?

VAEHVAEH:

yes please

Imagine:

I need this answered as well tho

VAEHVAEH:

they already helped right?

supie:

@Mercury?

VAEHVAEH:

think what we have to do is use this a+b+c=4 and a^2+b^2+c^2=10 and then a^3+b^3+c^3=22 To find what variable a, b, and c are

supie:

Thats what I said.

VAEHVAEH:

thats wat supie said

Imagine:

Still isn't solved tho

Imagine:

Smh

VAEHVAEH:

we cant tell the anwsers just help

Imagine:

smh

Imagine:

They can give you an answer if they back it up on why it's correct.

VAEHVAEH:

-_-

gabrielchestnut:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Imagine They can give you an answer if they back it up on why it's correct. \(\color{#0cbb34}{\text{End of Quote}}\) you deserve a badge for that

Imagine:

Oop- Thanks.

Imagine:

Bro I stumped Jhonny

Imagine:

Smh

darkknight:

a+b+c=4 and a^2+b^2+c^2=10 and then a^3+b^3+c^3=22 Then what does a^4+b^4+c^4=? a+b+c=4 a = 4-b-c (4-b-c)^2+b^2+c^2=10 \[b^2+2bc+c^2-8b-8c+16+b^2+c^2=10\] \[2b^2+2bc+2c^2-8b-8c+6=0\] \[2(b^2+bc+c^2-4b-4c+3)=0\] We also know that a^3+b^3+c^3=22 and a = 4-b-c so we rewrite this \[(4-b-c)^3+b^3+c^3=22\] simplify: \[(-b^3-3b^2c-3bc^2-c^3+12b^2+24bc+12c^2-48b-48c+64)+ b^3 + c^2 = 22\] \[(-3b^2c-3bc^2+12b^2+24bc+12c^2-48b-48c+64) = 22\]

Imagine:

NO WAY DID YOU JUST SOLVE THIS

darkknight:

So... \[2(b^2+bc+c^2-4b-4c+3)=0\] \[(-3b^2c-3bc^2+12b^2+24bc+12c^2-48b-48c+64) = 22\] 2 equations and 2 variables, from here you can solve for either b or c, and then plug that number back in, after that once you get b and c plug back into the first equation and solve for a

Imagine:

brooo

Imagine:

Not even jhonny solved this

Imagine:

Brooo

darkknight:

Once you get a, b, c then plug in values into a^4+b^4+c^4=?

darkknight:

Except one problem, how in the world do we find b and c? I made the equations but like...

darkknight:

what is this problem? no way i am doing geometry rn

Imagine:

Nope.

Imagine:

This is Advanced Calc.

Imagine:

._.

darkknight:

hmm, I think I got those equations correct but i don't know how to solve for b and c because of all the variables having different squares and such, it might be doable but it will def require a lot of math. These are just my thoughts on this but I can't solve the whole thing without committing hours to this.

Imagine:

That's good enough for me, I'll fiqure it out from here. Thanks

Imagine:

<3

darkknight:

np